# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ 3.2 Asymptotic properties and efficiency considerations
Given an estimator $\widehat{P}_{n}\in\mathcal{P}_{H}$ of $P_{0}$, Theorem 1 shows that the influence function $\varphi_{\widehat{P}_{n},H}$ characterizes the bias of the plug-in estimator, $\Psi(\widehat{P}_{n})-\Psi(P_{0})$, up to second-order remainder terms. To correct for this bias, the proposed DRL estimator $\psi_{n,H}$ augments the plug-in estimator with the empirical mean of the influence function. In a more general setting, we will show in Section 3.3 that $\psi_{n,H}$ is asymptotically linear for $\psi_{0}$, with influence function $\varphi_{0,H}$, under conditions. Consequently, $\psi_{n,H}-\psi_{0}=P_{n}\varphi_{0,H}+o_{p}(n^{-1/2})$, and $\sqrt{n}\{\psi_{n,H}-\psi_{0}\}$ converges in distribution to a $\mathcal{N}(0,\sigma_{0}^{2})$ random variable with variance $\sigma_{0}^{2}:=E_{0}[\{\varphi_{0,H}(S_{0},A_{0},Y_{0},S_{1})\}^{2}]$. We further show that the DRL estimator is rate doubly robust, meaning that asymptotic linearity holds provided the nuisance estimators are consistent and satisfy the rate condition:
$\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H}-q_{0})\|_{P_{0}}=o_{p}(n^{-1/2}).$
The limiting variance of $\psi_{n,H}$ equals the variance of its influence function, $\operatorname{Var}(\varphi_{0,H}(S_{0},A_{0},Y_{0},S_{1}))$. In the nonparametric case where $H=L^{\infty}(\lambda)$, Theorem 1 recovers the nonparametric efficient influence function (EIF) for the policy value $E_{0}[V^{\pi}(q_{0})(S_{0})]$ derived in Kallus and Uehara (2020, 2022). In this setting, $\psi_{n,H}$ is asymptotically efficient. However, under semiparametric restrictions, the influence function $\varphi_{0,H}$ in Theorem 1 generally differs from the EIF of $\Psi:\mathcal{P}_{H}\rightarrow\mathbb{R}$. As a result, $\psi_{n,H}$ may be inefficient under the model $\mathcal{P}_{H}$, though necessarily more efficient than fully nonparametric estimators under semiparametric restrictions. We note that efficient estimation under $\mathcal{P}_{H}$ often requires inverse weighting by the conditional variance of the outcome (Laan and Robins, 2003). Thus, while potentially inefficient, $\psi_{n,H}$ is often more stable and simpler to implement—much like ordinary least squares, which remains widely used despite its inefficiency under heteroscedasticity.
The limiting variance of $\psi_{n,H}$, and its potential efficiency gains over nonparametric estimators, are primarily governed by the variability of $\mathcal{T}_{0}(\alpha_{0,H})$, which reflects the degree of policy overlap and intertemporal state overlap. In the nonparametric case and for the policy value estimand, $\mathcal{T}_{0}(\alpha_{0,H})$ corresponds to the state occupancy ratio $d_{0}$ defined in (2), which may exhibit high variance or fail to exist when overlap is limited (Mehrabi and Wager, 2024). Semiparametric restrictions on the $Q$-function generally reduce the variability of $\mathcal{T}_{0}(\alpha_{0,H})$, becoming less sensitive to limited overlap. Specifically, $\mathcal{T}_{0}(\alpha_{0,H})$ corresponds to the $L^{2}(P_{0,A_{0},S_{0}})$-projection of its nonparametric counterpart onto the range $\mathcal{T}_{0}(\overline{H}_{P_{0}})$, assuming the latter exists. Since projection contracts norms, the variance of $\varphi_{0,H}$ typically decreases as the function class $H$ becomes more restrictive. For example, if $H$ constrains $q_{0}$ to depend only on a subvector $\widetilde{S}_{t}\subset S_{t}$, then $\mathcal{T}_{0}(\alpha_{0,H})(A_{0},S_{0})$ reduces to the conditional expectation of $d_{0}(A_{0},S_{0})$ given $(A_{0},\widetilde{S}_{0})$, requiring overlap only within this lower-dimensional subspace. These efficiency gains can be substantial, as the following example illustrates in the context of data fusion.
Example 3 (Long-term causal effect in a data fusion setting).
Let $A_{t}:=Z$ be a time-homogeneous study assignment, and define the state as $S_{t}=(Z,\widetilde{S}_{t})$, as in Example 2, with $P_{0}(Z=1\mid S_{0})=p$. We consider a data fusion setting where experimental data ($Z=1$) is augmented with historical controls ($Z=0$), extending the approach of Kallus et al. (2018) and van der Laan et al. (2024d) to MDPs. The expected cumulative reward in the treatment arm, $\mathbb{E}_{0}[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}\mid Z=1]$, is identified by $E_{0}[q_{0}^{\pi_{0}}(1,(1,\widetilde{S}_{0}))\mid Z=1]$, where $\pi_{0}(a\mid(z,\widetilde{s})):=\mathbbm{1}(a=z)$ is the behavior policy. Assume the historical control dataset is large, so the control $Q$-function $V^{\pi}_{\text{hist}}(\widetilde{s}):=q_{0}(0,(0,\widetilde{s}))$ is effectively known. Consider the offset model $H:=\{q:(a,(z,\widetilde{s}))\mapsto V^{\pi}_{\text{hist}}(\widetilde{s})+\kappa+\beta z;\kappa,\beta\in\mathbb{R}\}$, which assumes $V^{\pi}_{\text{hist}}$ is known and posits that the data-combination bias $q_{0}(1,(1,\widetilde{S}_{0}))-q_{0}(0,(0,\widetilde{S}_{0}))$ is constant in $\widetilde{S}_{0}$. Although $H$ is affine, Theorem 2 still applies, with $\alpha_{0,H}$ replaced by $\alpha_{0,T_{0}(H)}$, where $T_{0}(H):=H-\{V^{\pi}_{\text{hist}}\}$ is the tangent space at $P_{0}$. Under standard MDP conditions, the nuisance term $\mathcal{T}_{0}(\alpha_{0,T_{0}(H)})$ in the influence function is $(a,(z,\widetilde{s}))\mapsto\frac{z}{p}\cdot E_{0}[d_{0}(Z,\widetilde{S}_{0})\mid Z=1]$, where $d_{0}(z,s)=\frac{d\mathbb{P}_{0}(\widetilde{S}_{t}=s\mid Z=z)}{d\mathbb{P}_{0}(\widetilde{S}_{0}=s\mid Z=z)}$. Since $(1-\gamma)E_{0}[d_{0}(Z,\widetilde{S}_{0})\mid Z=z]=1$ almost surely, the influence function—and thus the asymptotic variance of $\psi_{n,H}$—is unaffected by the degree of intertemporal state overlap. ∎
3.3 Extension to model-robust inference and asymptotic theory
In practice, the $Q$-function $q_{0}$ may not lie exactly in the model $H\subset L^{\infty}(\lambda)$, but may be well-approximated by elements of $H$, with potential efficiency gains justifying misspecification bias (Crump et al., 2006; van der Laan et al., 2023). Alternatively, $H$ may serve as a working model, with interest focused on a linear functional of the best approximation to $q_{0}$, such as the best linear predictor when $H$ is linear (Whitney et al., 2020; Vansteelandt and Dukes, 2022; Chambaz et al., 2012; Chernozhukov et al., 2018b). To enable valid inference without assuming $P_{0}\in\mathcal{P}_{H}$, we define a projection parameter $\Psi_{H}:\mathcal{P}\rightarrow\mathbb{R}$ that extends $\Psi:\mathcal{P}_{H}\rightarrow\mathbb{R}$ to the full nonparametric model. That is, $\Psi_{H}$ agrees with $\Psi$ on the submodel $\mathcal{P}_{H}$, but remains well-defined and interpretable under misspecification (Buja et al., 2019). We show how the DRL estimator $\psi_{n,H}$ can be adjusted to target $\Psi_{H}$, enabling model-robust inference. We then establish the asymptotic linearity and efficiency of both $\psi_{n,H}$ and its bias-corrected version as estimators of $\Psi_{H}(P_{0})$ under correct and incorrect specification, respectively.
We consider inference on the projection estimand $\psi_{0,H}:=E_{0}[m(S_{0},A_{0},q_{0,H})]$, where $q_{0,H}$ is a projection of the $Q$-function onto the model $H$. For each $P\in\mathcal{P}$, we define the Bellman projection $q_{P,H}$ as a solution to
$q_{P,H}\in\arg\min_{q\in\overline{H}_{P}}E_{P}\left[\left\{Y_{0}-\mathcal{T}_{P}(q)(A_{0},S_{0})\right\}^{2}\right],$ (8)
which exists and is unique under the invertibility condition on $\mathcal{T}_{P}$ in Condition C1(ii). The resulting working parameter $\Psi_{H}:\mathcal{P}\to\mathbb{R}$, defined by $\Psi_{H}(P):=E_{P}[m(S_{0},A_{0},q_{P,H})]$, extends $\Psi$ to the nonparametric model $\mathcal{P}$ while preserving its value on $\mathcal{P}_{H}$. Moreover, $q_{P,H}$ satisfies the projected Bellman equation $\mathcal{T}_{P}(q_{P,H})=\mu_{P,H}$, where $\mu_{P,H}:=\arg\min_{\mu\in\mathcal{T}_{P}(\overline{H}_{P})}\|\mu_{P}-\mu\|_{P}$ is the $L^{2}(P)$-projection of the outcome regression $\mu_{P}$ onto the range $\mathcal{T}_{P}(\overline{H}_{P})$. Hence, $q_{P,H}$ is the best approximation to $q_{P}$ within $H$ under the norm $\|\mathcal{T}_{P}(\cdot)\|_{P}$.
The following theorem extends the pathwise differentiability result for $\Psi$ in Theorem 1 to the projection parameter $\Psi_{H}$. In the following theorem, for each distribution $P\in\mathcal{P}$, we define the influence function $\varphi^{*}_{P,H}$ as the map:
$(s,a,y,s^{\prime})$ $\mapsto\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V^{\pi}(q_{P,H})(s^{\prime})-q_{P,H}(a,s)\}$
+$\quad+\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(s^{\prime})-\mathcal{T}_{P}(\alpha_{P,H})(a,s)\}\{\mu_{0}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}$
$\quad+m(s,a,q_{P,H})-\Psi_{H}(P).$
Theorem 2 (Pathwise differentiability and efficient influence function).
@@ 3.3 Extension to model-robust inference and asymptotic theory
Suppose C1 holds. Then the parameter $\Psi_{H}:\mathcal{P}\rightarrow\mathbb{R}$ is pathwise differentiable at $P_{0}$, with efficient influence function $\varphi^{*}_{0,H}$. Moreover, for any $\overline{P}\in\mathcal{P}$ for which $\varphi^{*}_{\overline{P},H}$ exist, the parameter satisfies the expansion: $\Psi_{H}(\overline{P})-\Psi_{H}(P_{0})=-P_{0}\varphi^{*}_{\overline{P},H}+R_{H}^{*}(\overline{P},P_{0}),$ where
@@
$\quad+\left\langle(\mathcal{T}_{\overline{P}}-\mathcal{T}_{0})(\alpha_{\overline{P},H}),\,\mu_{0}-\mu_{\overline{P}}+\mathcal{T}_{\overline{P}}(q_{\overline{P},H})-\mathcal{T}_{0}(q_{0,H})\right\rangle_{P_{0}}.$
When $q_{0}\in H$, the EIF $\varphi^{*}_{0,H}$ in Theorem 2 reduces to the influence function $\varphi_{0,H}$ from Theorem 1 for the restricted parameter $\Psi\colon\mathcal{P}_{H}\rightarrow\mathbb{R}$. Under misspecification of $H$, $\varphi^{*}_{0,H}$ includes additional terms involving the approximation error $\mathcal{T}_{0}(q_{0}-q_{0,H})$ and the residual $(s,a,s^{\prime})\mapsto\alpha_{0,H}(a,s)-\gamma V^{\pi}(\alpha_{0,H})(s^{\prime})-\mathcal{T}_{0}(\alpha_{0,H})(a,s).$ The remainder in the von Mises expansion also includes additional terms involving the errors $\mathcal{T}_{P}-\mathcal{T}_{0}$ and $\mu_{P}-\mu_{0}$.
The model-robust EIF of Theorem 2 demonstrates that, to correct the bias of the DRL estimator $\psi_{n,H}$ for the projection estimand $\Psi_{H}(P_{0})$ under model misspecification, it suffices to add an additional bias correction term. In particular, a model-robust estimator of $\Psi_{H}(P_{0})$ is given by
$\psi_{n,H}^{*}$ $=\psi_{n,H}+\frac{1}{n}\sum_{i=1}^{n}\big{(}\mu_{n}-\widehat{\mathcal{T}}_{n}(q_{n,H})\big{)}(A_{0,i},S_{0,i})\left\{\alpha_{n,H}(A_{0,i},S_{0,i})-\gamma V^{\pi}(\alpha_{n,H})(S_{1,i})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\right\}.$
where $q_{n,H}\in H$, $\alpha_{n,H}\in H$, $\widehat{\mathcal{T}}_{n}(q_{n,H})$, $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$, and $\mu_{n}$ are estimators of $q^{\pi}_{0,H}$, $\alpha_{0,H}$, $\mathcal{T}_{0}(q_{0,H})$, $\mathcal{T}_{0}(\alpha_{0,H})$, and $\mu_{0}$, respectively. The estimator $\psi_{n,H}^{*}$ generalizes the DRL estimator $\psi_{n,H}$ from the previous section, which is recovered under correct specification by setting $\mu_{n}:=\widehat{\mathcal{T}}_{n}(q_{n,H})$. Unlike $\psi_{n,H}$, however, the model-robust estimator $\psi_{n,H}^{*}$ enables efficient inference for $\Psi_{H}(P_{0})$ regardless of whether $H$ is correctly specified.
In the following conditions and theorem, let $\varphi_{n,H}^{*}$ denote the estimator of the EIF $\varphi_{0}^{*}$ from Theorem 2, obtained by plugging in our nuisance estimators.
@@
    Consistency: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\}=o_{p}(1)$.
@@
      $\left\|\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\right\|_{P_{0}}\cdot\left\|\widehat{\mathcal{T}}_{n}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})+\mu_{0}-\mu_{n}\right\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
@@
Assume C1, C2, and C3. Then, $\psi_{n,H}^{*}-\psi_{0,H}=(P_{n}-P_{0})\varphi_{0,H}^{*}+o_{p}(n^{-\frac{1}{2}})$. Moreover, $\psi_{n,H}^{*}$ is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model .
@@
Condition C2 is an empirical process condition that holds if $\|\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\|_{P_{0}}=o_{p}(1)$ and the difference lies in a Donsker class, or if sample-splitting or cross-fitting is used to bypass Donsker conditions (van der Laan et al., 2011; Chernozhukov et al., 2018a). Condition C3 is a doubly robust rate condition requiring that the nuisance estimators converge at sufficiently fast rates. The first rate condition, $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\|_{P_{0}}=o_{p}(n^{-1/2})$, commonly appears in nonparametric inference for functionals of solutions to inverse problems (Kallus and Uehara, 2022; Bennett et al., 2023a; Li et al., 2024). Under C1, this condition holds if both $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$ and $q_{n,H}$ converge to their targets in $\|\cdot\|_{P_{0}}$-norm at rates faster than $n^{-1/4}$. Under correct specification ($q_{0}\in H$), the second rate condition is trivially satisfied when using the DRL estimator $\psi_{n,H}$, since $\mu_{n}=\widehat{\mathcal{T}}_{n}(q_{n,H})$ and $\mu_{0}=\mathcal{T}_{0}(q_{0,H})$. These conditions can be satisfied under appropriate smoothness assumptions using machine learning algorithms such as neural networks, random forests, and gradient-boosted trees.
3.4 Model misspecification bias and data-driven model selection
Selecting an appropriate working model $H$ for $q_{0}$ is challenging and may compromise inference due to misspecification bias. While choosing a rich model $H$ may mitigate such bias, it comes at the cost of increased variance and greater reliance on intertemporal overlap conditions. To navigate this bias-variance trade-off, it is natural to estimate the functional form of the $Q$-function $q_{0}$ from data. However, conventional wisdom holds that such data-driven choices can invalidate inference (Leeb and Pötscher, 2005). In this section, we propose an adaptive estimator (van der Laan et al., 2023) that learns a data-dependent model $H_{n}$ and establish conditions for valid inference using sieve methods and model selection. To do so, we show that misspecification of the $Q$-function induces only second-order bias in the target functional.
Let $H_{n}\subseteq H$ be a data-driven working model for the $Q$-function $q_{0}$. For example, $H_{n}$ may be selected via cross-validated FQI over a sieve of models $H_{1}\subset H_{2}\subset\cdots\subset H_{\infty}:=H$, where $H$ is correctly specified and contains $q_{0}$. Alternatively, $H_{n}$ could result from variable selection or a learned feature transformation (Pritz et al., 2021; Pavse and Hanna, 2024). Given estimators $q_{n,H_{n}}\in H_{n}$ and $\alpha_{n,H_{n}}\in H_{n}$ of $q_{0,H_{n}}$ and $\alpha_{0,H_{n}}$, our proposed adaptive DRL estimator of $\Psi(P_{0})$ takes the form of a one-step estimator that naively assumes $H_{n}$:
$\psi_{n,H_{n}}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H_{n}})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H_{n}})(S_{1,i})-q_{n,H_{n}}(A_{0,i},S_{0,i})\right\}.$
To analyze its asymptotic behavior, we assume $H_{n}$ approximates a fixed but unknown oracle submodel $H_{0}\subseteq H$ containing $q_{0}$. Provided the approximation error between $H_{n}$ and $H_{0}$ vanishes suitably, the Adaptive Debiased Machine Learning (ADML) framework of van der Laan et al. (2023) implies that $\psi_{n,H_{n}}$ is asymptotically equivalent to the oracle estimator $\psi_{n,H_{0}}$, which assumes knowledge of $H_{0}$. We formally prove this result in Appendix B.
Our novel contribution is the following theorem, which shows that the parameter approximation bias $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ is second order in the model approximation error and thus asymptotically negligible under suitable conditions. This result extends related bounds for linear functionals of the outcome regression derived in van der Laan et al. (2023) to the MDP and inverse problem setting.
Theorem 4 (Second-order model approximation error).
where $H_{n,0}:=H_{n}\oplus H_{0}$ is the direct sum linear model.
A consequence of Theorem 4 is that the prespecified DML estimators $\psi_{n,H}$ and $\psi_{n,H}^{*}$ from the previous sections are robust to mild misspecification of the model $H$. Hence, these estimators may still provide valid inference for $\psi_{0}$, provided that the bias due to misspecification is negligible relative to the standard error of the estimator. In Appendix B, we show that the adaptive DRL estimator satisfies $\psi_{n,H_{n}}=\psi_{n,H_{0}}+o_{p}(n^{-1/2})$, provided that $\|\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}})\|\cdot\|\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\|_{P_{0}}=o_{p}(n^{-1/2})$, and that the data-dependent influence function $\varphi_{0,H_{n}}$ converges in probability to the oracle influence function $\varphi_{0,H_{0}}$. This latter condition imposes a mild requirement on the consistency and stability of the model selection procedure but can be avoided if the model is learned using data independent of that used to evaluate the estimator (van der Laan et al., 2023).
For the approximation error $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ to vanish, both $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ must converge to one another in $L^{2}(P_{0})$, and $\mathcal{T}_{0}(q_{0,H_{n}})$ must converge to $\mathcal{T}_{0}(q_{0})$. This convergence requires that the learned model $H_{n}$ approximates both the true $Q$-function $q_{0}$ and the union model representer $\alpha_{0,H_{n,0}}$ with vanishing error in the $\|\mathcal{T}_{0}(\cdot)\|$ norm. In sieve-based model selection, the event $H_{n}\subseteq H_{0}$ typically holds with high probability, in which case $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})=\mathcal{T}_{0}(\alpha_{0,H_{0}})$, and it suffices for $H_{n}$ to grow sufficiently fast. For general model selection procedures, convergence of $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ to $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ further requires that any directions in $H_{n,0}\cap H_{n}^{\perp}$ contribute negligibly to the union model representer $\alpha_{0,H_{n,0}}$.
To further clarify these conditions, suppose the working model $H_{n}:=H_{\phi_{n}}$ and the oracle model $H_{0}:=H_{\phi_{0}}$ are induced by feature transformations. For a transformation $\phi:\mathcal{A}\times\mathcal{Z}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}$. The combined model $H_{n,0}$ is given by $H_{(\phi_{n},\phi_{0})}$, where $(\phi_{n},\phi_{0})$ denotes the feature map formed by stacking $\phi_{n}$ and $\phi_{0}$. Theorem 4 implies that the approximation bias vanishes if the nuisance functions derived from $\phi_{n}$ and $(\phi_{n},\phi_{0})$ converge to those derived from $\phi_{0}$. In Lemma 14 of Appendix H.1, we show that with features $X$ and outcome $Y$, the mean squared error between $E_{0}[Y\mid\phi_{0}(X)]$ and $E_{0}[Y\mid\phi_{n}(X),\mathcal{D}_{n}]$ is bounded by the feature approximation error $\int\|\phi_{n}(x)-\phi_{0}(x)\|_{\mathbb{R}^{m}}^{2}\,P_{0,X}(dx)$. A sufficient condition for this bound is that the map $(t_{1},t_{2})\mapsto E_{0}[Y\mid\phi_{n}(X)=t_{1},\phi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely Lipschitz continuous.
We revisit this setup in Section 5, where we introduce an adaptive DRL estimator that leverages $Q$-function estimates as a data-adaptive form of dimension reduction.
4 Methods and challenges in nuisance estimation
In this section, we describe how the $Q$-function $q_{0,H}$ and the Riesz representer $\alpha_{0,H}$ can be estimated using flexible machine learning methods. The $Q$-function can be readily estimated using standard reinforcement learning methods. In contrast, estimating the Riesz representer is more challenging, as it involves solving a convex–concave min–max problem. Recognizing this challenge, the next section proposes a novel DRL estimator that achieves debiasing without requiring estimation of the Riesz representer.
A popular algorithm for estimation of $q_{0,H}$ in reinforcement learning is fitted Q-iteration (FQI) (Munos and Szepesvári, 2008), an iterative method for solving the Bellman integral equation (1). FQI is based on the observation that if $q_{0,H}$ were known, it could be estimated by regressing $Y_{0}+\gamma V^{\pi}(q_{0,H})(S_{1})$ on $(A_{0},S_{0})$, since rearranging (1) yields $E_{0}[Y_{0}+\gamma V^{\pi}(q_{0,H})(S_{1})\mid A_{0},S_{0}]=q_{0,H}(A_{0},S_{0})$ almost surely. Since $q_{0,H}$ is unknown, FQI initializes $q_{n,H}^{\pi,(0)}:=0$ and iteratively updates it by regressing the Bellman outcome $Y_{0}+\gamma V^{\pi}(q_{n,H}^{\pi,(k)})(S_{1})$ on $(A_{0},S_{0})$ over the model class $H$ at each iteration $k+1\in\mathbb{N}$. Iteration stops when the $\ell^{2}$ norm between consecutive updates is sufficiently small or when out-of-sample or cross-validated risk ceases to improve. Algorithm 1 details the procedure. Theoretical guarantees on validity and convergence rates of FQI using generic function approximation methods, such as neural networks (Bishop, 1994), random forests (Breiman, 2001), and gradient boosted trees (Friedman, 2001), are provided in Munos and Szepesvári (2008) and Agarwal et al. (2019).
Algorithm 1 Fitted Q-Iteration
0:  Data, Function class $H$, number of iterations $K$;
1:  Initialize $q_{n,H}^{\pi,(0)}:=0$;
2:  for $k=0,1,\ldots,K-1$ do

### v4

@@ 3.2 Model-robust extension and asymptotic theory
Suppose C1 holds. Then the parameter $\Psi_{H}:\mathcal{P}\to\mathbb{R}$ is pathwise differentiable at $P_{0}$ with efficient influence function $\varphi^{*}_{0,H}$. Moreover, for any $\overline{P}\in\mathcal{P}$ for which $\varphi^{*}_{\overline{P},H}$ exists, the parameter admits the expansion
$\Psi_{H}(\overline{P})-\Psi_{H}(P_{0})$ $=-\,P_{0}\varphi^{*}_{\overline{P},H}+R_{H}^{*}(\overline{P},P_{0}),$
@@
$\qquad+\left\langle(\mathcal{T}_{\overline{P}}-\mathcal{T}_{0})(\alpha_{\overline{P},H}),\,r_{0}-r_{\overline{P}}+\mathcal{T}_{\overline{P}}(q_{\overline{P},H})-\mathcal{T}_{0}(q_{0,H})\right\rangle_{P_{0}}.$
The variance of the EIF $\varphi^{*}_{P_{0},H}$ determines the $P_{0}$-generalized Cramér–Rao efficiency bound for the projection parameter $\Psi_{H}$. When semiparametric restrictions are imposed through $H$, this efficiency bound can be strictly smaller than that of the unrestricted parameter $\Psi$; we return to this point in the next section. Specializing Theorem 1 to the case $H=L^{\infty}(\lambda)$ and to the policy value functional recovers the nonparametric setting of kallus2020double; kallus2022efficiently. Under correct specification ($P_{0},\overline{P}\in\mathcal{P}_{H}$), the second term in the EIF $\varphi^{*}_{\overline{P},H}$ and the second term in the remainder $R_{H}^{*}(\overline{P},P_{0})$ both vanish.
The following theorem establishes double robustness, asymptotic normality and efficiency of $\psi_{n,H}^{*}$ for $\Psi_{H}$. Let $\varphi_{n,H}^{*}$ denote the plug-in estimator of the efficient influence function $\varphi_{0,H}^{*}$ from Theorem 1, obtained by substituting the estimated nuisance functions.
@@
    Negligible empirical process: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\}=o_{p}(1)$.
@@
      $\left\|\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\right\|_{P_{0}}\cdot\left\|\widehat{\mathcal{T}}_{n}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})+r_{0}-r_{n}\right\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
@@
Assume C1, C2, and C3. Then, $\psi_{n,H}^{*}-\psi_{0,H}=(P_{n}-P_{0})\varphi_{0,H}^{*}+o_{p}(n^{-\frac{1}{2}})$. Moreover, $\psi_{n,H}^{*}$ is a $P_{0}$-regular and efficient estimator for $\Psi_{H}$ under the nonparametric model .
@@
Condition C2 is an empirical process requirement that holds if $\|\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\|_{P_{0}}=o_{p}(1)$ with the difference lying in a Donsker class, or if sample-splitting or cross-fitting is used to bypass Donsker conditions (van2011cross; chernozhukov2018double). Condition C3 is a doubly robust rate condition requiring sufficiently fast convergence of the nuisance estimators. The first rate condition, $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\|_{P_{0}}=o_{p}(n^{-1/2})$, generalizes the standard requirement for nonparametric inference on the policy value (kallus2022efficiently; bennett2023minimax; li2024regularized). The second condition is specific to the semiparametric setting and applies only to the model-robust estimator $\psi_{n,H}^{*}$. Under correct specification ($q_{0}\in H$), it is automatically satisfied by the uncorrected estimator $\psi_{n,H}$, which uses $r_{n}:=\widehat{\mathcal{T}}_{n}(q_{n,H})$ and $r_{0}:=\mathcal{T}_{0}(q_{0,H})$.
3.3 Efficiency considerations.
The limiting variance of the semiparametric DRL estimator $\psi_{n,H}^{*}$ is determined by the variability of the weighting function $\mathcal{T}_{0}(\alpha_{0,H})$. For the policy value, this weighting function equals the state–occupancy ratio $d_{0}$ in the fully nonparametric case, and otherwise its $L^{2}(P_{0})$-projection onto the range $\mathcal{T}_{0}(\overline{H}_{P_{0}})$ (assuming $d_{0}$ has finite variance). Thus, the semiparametric efficiency bound is strictly smaller whenever $H$ excludes the nonparametric Riesz representer $\alpha_{0,L^{\infty}(\lambda)}$. For example, if $H$ consists of functions linear in a feature map $\varphi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{k}$, then the weighting function becomes the best linear predictor of $d_{0}$, namely $\mathcal{T}_{0}(\alpha_{0,H})(a,s)=\beta_{0}^{\top}\varphi_{0,\gamma}(a,s)$, where $\beta_{0}:=\arg\min_{\beta}E_{0}[(d_{0}(A_{0},S_{0})-\beta^{\top}\varphi_{0,\gamma}(A_{0},S_{0}))^{2}]$ and $\varphi_{0,\gamma}(a,s):=\varphi(a,s)-\gamma\,E_{0}[\int\varphi(a^{\prime},S_{1})\,\pi(a^{\prime}\mid S_{1})\,\nu(da^{\prime})\mid A_{0}=a,S_{0}=s]$, with expectations of vectors taken component-wise. Efficiency gains are largest when the high-variance components of $d_{0}$ lie outside the linear span of $\varphi_{0,\gamma}$.
Similar efficiency gains have been established in linear MDPs, where both the transition kernel and the reward function are assumed to be linear in a known feature map (xie2023semiparametrically). In contrast, we obtain the same benefits under substantially weaker assumptions: it is enough to restrict only the $Q$-function, without imposing a parametric model on the underlying MDP. Intuitively, constraining the $Q$-function implicitly restricts both the reward and the transition kernel through the Bellman equation. The following example makes this efficiency phenomenon explicit in a simple illustrative setting.
Example 3 (Eliminating overlap dependence via time-invariant state structure).